OpenStudy (anonymous):

John began his job making $25 the first day. After that he was paid $6.75 per hour. Write the linear equation for this problem in y=mx+b form.

OpenStudy (frostbite):

y= 6.75 * x + 25 Money earned = $6.75 per hour * time + $25 Simple relation.

OpenStudy (anonymous):

Thank you so much! you earned yourself a fan :D

OpenStudy (frostbite):

Think of it as this way: every time you see a unit with fx. money/hour, cats/time, dogs/train.... you know that is your m value becuase m is defined as delta Y/Delta X so it is the y unit divided with the x unit

OpenStudy (frostbite):

But no problem

OpenStudy (anonymous):

Can you answer some other questions i have?

OpenStudy (frostbite):

Sure I need to go to a lecture ^^

OpenStudy (frostbite):

But good luck.

OpenStudy (anonymous):

Ohh its okay then. I became a fan. thanks alot :)

OpenStudy (anonymous):

Very easy

OpenStudy (anonymous):

The y=mx+b is a pretty simple equation once you understand it. The Y is going to be the answer once the equation is solved. The X is the variable that changes. So for 2, the values are 0, 20 and 40.

OpenStudy (anonymous):

The M in this problem is what the X is multiplied by, so in this problem, it is 6.75. y=6.75x+25

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

yeah i get it now! xD i sent you another question!

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

how do i give a medal?

OpenStudy (anonymous):

Np:)

OpenStudy (anonymous):

I'll give you one for that understanding

OpenStudy (anonymous):

thanks :) but i have more questions. :(